OpenStudy (anonymous):

A store mixes Kenyan Coffee worth $11 per kilogram and Turkish coffee worth $15 per kilogram. The mixture is to sell for $14 per kilogram. Find how much of each should be used to make a 480-kilogram mixture.How many Kilograms of the Kenyan coffee should be in the mixture?_kgHow many Kilograms of the Turkish coffee should be in the mixture?_kg i cant figure this out.

OpenStudy (anonymous):

360 kilograms of Turkish coffe and 120 kilograms of Kenyan coffee

OpenStudy (anonymous):

Because: x+y=1 11x+15y=14(1) 11x-15(1-x)=14 -4=-1 x=(1/4) So, for every kilogram, 1/4 should be kenyan coffee. Since there are 480 kilograms, multiply 480 by 1/4 to get 120. The rest must be Turkish coffee, to 480-120=360, the amount of turkish coffee

OpenStudy (anonymous):

Does that make sense?